# A Lost Opportunity for Vision-Language Models: A Comparative Study of Online Test-time Adaptation for Vision-Language Models

### v1

@@ 4 Updating Model Parameters with Test-time Adaptation
While the focus of the previous section has been on leveraging prompts and vision-space ensembles, in this section we want to put emphasis on a surprisingly underexplored topic, namely leveraging existing TTA methods for adapting vision-language models. The idea is straightforward, the text encoder of a CLIP model is frozen, allowing to precompute the text embeddings $\{\bm{t}_{1},\bm{t}_{2},\dots,\bm{t}_{K}\}$, ensuring efficiency during prediction. Given an image embedding $\bm{z}_{t}$, the cosine similarity can be computed $s_{t}=\mathrm{sim}(\bm{t}_{k},\bm{z}_{t})$. Treating the cosine similarities as the network’s logits, the output probabilites can be received through Eq. 1. In this way we can treat any CLIP model as a common image classifier, enabling the application of any existing TTA method for image classification. Note that it is also possible to update the text encoder’s parameters, but for now, we limit our analysis to only updating the parameters of the image encoder. In the following experiments a batch size of 64 is employed.
4.1 Test-time Normalization for CLIP
First, we investigate the performance of BN–1 which recalculates the batch normalization (BN) statistics using the current test batch. While Schneider et al. [41] showed that recalculating the batch normalization statistics during test-time can significantly reduce the error rate for models pretrained on ImageNet, we investigate whether this is also the case for a CLIP model that was trained on millions of data samples covering a much broader data distribution. In Figure 3 the zero-shot performance (source) and BN–1 performance is illustrated for CLIP with a RN50 and RN101 backbone using a single prompt. It can be clearly seen that unlike for models pretrained on ImageNet, the average performance across the investigated datasets substantially decreases when applying BN–1. For RN50 the average error rate increases from 50.2% to 74.1% and for RN101 from 46.8% to 71.0%. This can be possibly attributed to much larger batch sizes and a much broader data distribution used during CLIP pretraining. A similar phenomenon is described in [29], where employing BN–1 for a regular ImageNet pretrained RN50 decreases the error rate on ImageNet-C from 82.0% to 68.6% in a continual TTA setting, but increases to 82.5% in a mixed-domains TTA setting, where all corruptions of ImageNet-C are randomly suffled within the test sequence. Since BN–1 is employed by most TTA methods during adaptation, we conclude that RN backbones are not feasible. Instead we focus our following analysis on vision transformer backbones that do not employ BN.
@@ 4.2 Are Existing TTA Methods Beneficial for Vision-language Foundation Models?
Figure 5: Comparison of different models sorted according to their number of model parameters from low (left) to high (right). The average error rate across all datasets is reported.
@@
Since the previously investigated continual TTA setting might not apply to real-world online data streams, we additionally investigate a scenario with temporally correlated samples. In the correlated TTA setting the data of each domain is sorted by the class label rather than randomly shuffled, resulting in class-imbalanced batches. The results are reported in Table 3. For the more challenging correlated TTA setting, in contrast to the continual setting, not all TTA methods are capable to improve upon the source performance. Only CMF and ROID show a stable adaptation. Due to the employed prior correction [^1] In contrast to the original prior correction, we find that applying the prior correction in the output probability space instead of the logit space shows a more consistent performance for the investigated CLIP models. proposed in [29], CMF and ROID even report a better performance than in the continual setting.
@@ 4.5 Limitations
A limitation of applying existing TTA methods to vision-language models is that they often require batches of data to perform updates. As discussed in [29] this is only partially true. Using networks that do not employ BN layers, such as VisionTransformer [10] allow to recover the batch TTA setting by simply accumulating the gradients of the last $b$ test samples before updating the model. This comes with no computational overhead and even significantly reduces the memory requirement.

### v2

@@ 4.2 Are Existing TTA Methods Beneficial for Vision-Language Foundation Models?
Figure 5: Comparison of different models sorted according to their number of parameters from low (left) to high (right). The average error rate across all datasets is reported.
@@
Since the previously investigated continual TTA setting might not apply to real-world online data streams, we additionally investigate a scenario with temporally correlated samples. In the correlated TTA setting the data of each domain is sorted by the class label rather than randomly shuffled, resulting in class-imbalanced batches. The results are reported in Table 3. For the more challenging correlated TTA setting, in contrast to the continual setting, not all TTA methods are capable to improve upon the source performance. Only CMF and ROID show a stable adaptation. Due to the employed prior correction [^1] In contrast to the original prior correction, we find that applying the prior correction in the output probability space instead of the logit space shows a more consistent performance for the investigated CLIP models. proposed in [29], CMF and ROID perform even better than in the continual setting.
@@ 4.5 Limitations
A limitation of applying existing TTA methods to vision-language models is that they often require a batch of test data at each time step $t$ for effective parameter updates. However, as discussed in [29], this is only partially true. Networks that do not employ BN layers, such as VisionTransformer [10], allow to recover the batch TTA setting by simply accumulating the gradients of the last $b$ test samples before updating the model. This comes with no computational overhead and even significantly reduces the memory requirement.